# Verifying Equilibria in Finite-Horizon Probabilistic Concurrent Game Systems

### v5

@@ Abstract.
Finite-horizon probabilistic multiagent concurrent game systems, also known as finite multiplayer stochastic games, are a well-studied model in computer science due to their ability to represent a wide range of real-world scenarios involving strategic interactions among agents over a finite amount of iterations (given by the finite-horizon). The analysis of these games typically focuses on evaluating (verifying) and computing (synthesizing/realizing) which strategy profiles (functions that represent the behavior of each agent) qualify as equilibria. The two most prominent equilibrium concepts are the Nash equilibrium and the subgame perfect equilibrium, with the latter considered a conceptual refinement of the former. However, computing these equilibria from scratch is often computationally infeasible. Therefore, recent attention has shifted to the verification problem, where a given strategy profile must be evaluated to determine whether it satisfies equilibrium conditions. In this paper, we demonstrate that the verification problem for subgame perfect equilibria lies in PSPACE, while for Nash equilibria, it is EXPTIME-complete. This is a highly counterintuitive result since subgame perfect equilibria are often seen as a strict strengthening of Nash equilibria and are intuitively seen as more complicated.
@@ 1. Introduction
The ubiquity of distributed systems in modern computing architectures has motivated much recent work into the study of multi-agent systems [GTW02, Gup16, MMPV14, SL09, VdHW08, Woo09, AKM+19, RBV23, RV22, RV21, BBM10, BBMU11, BMS14, BBMU15, BRRvdB24, BRvdB23], which are abstract mathematical models in which two or more rational agents interact with each other in order to achieve their goals. The broad motivation behind these works is to provide a mathematical framework for analyzing the overall trends of the complex mechanics that arise when multiple agents interact with one another in a strategic manner.
@@
Much of the relevant work in this area comes from the field of formal methods, which has recently begun to analyze equilibria in multi-agent systems as a way to analyze “stable” behaviors in multi-agent systems, and, therefore, analyze the system’s properties as a whole. The two major problems in the field of formal methods are the realizability [PR89] and the verification problems [Ros92, RV22]. Therefore, much of the work in formal analysis of multi-agent systems is concerned with determining the existence of (the realizability problem) and verifying (the verification problem) different types of equilibria [GTW02, BBMU11, BBM10, BBMU15, RBV23, RV22, RV21, Gup16, GNPW18, GPW17, BBMU15, BRRvdB24, BRvdB23]. As evidenced by the previously cited works, however, work in formal verification often comes with a strong preference for deterministic systems and deterministic agent behaviors. This creates something of an ideological divide between equilibrium analysis in the formal methods literature and the study of equilibria in the game theory literature. Our broad goal in this paper is to begin to bridge this divide by applying this formal methods style of equilibrium analysis to probabilistic systems. This is a new line of inquiry, as even papers in the formal methods literature that consider probabilistic systems such as [GHL+21] (which itself notes the focus on deterministic systems in previous literature) consider notions of almost-sure satisfaction on infinite-length executions of a system. These notions allow for these probabilistic systems to be analyzed through the same techniques used for deterministic ones.
@@
In order to differentiate our work from this type of analysis, we consider finite-horizon systems, meaning that the system has an explicitly given finite bound on the number of steps in the execution of the system. This terminology follows the notation from the literature on Markov Decision Processes [Put94, Ete21] and breaks from the terminology of the formal verification literature, where finite-horizon is usually shorthand for “finite but unbounded” [Woo09, GV13]. Finite-horizon systems are also arguably more popular than their infinite-horizon counterparts in the literature on Markov Decision Process [Put94, Ete21, Sol22, BKN+19], which can be seen as the “single-agent” analog of the systems we consider in this paper.
@@
We wish to consider the computational complexity of the realizability and verification problems in finite-horizon probabilistic multi-agent systems. The two most popular equilibrium concepts in the field of game theory are the Nash equilibrium [Nas50] and the subgame perfect equilibrium [OR94], with the latter a strengthening of the former. However, this immediately presents a problem for the realizability problem. Since the realizability problem is concerned with determining whether an equilibrium exists and we are considering probabilistic systems with a finite-horizon, the realizability problem becomes trivial. This is because subgame perfect equilibria can be shown to always exist in these systems through the use of algorithms like backwards induction [Far24]. Furthermore, if we note that every subgame perfect equilibrium is, by definition, a Nash equilibrium, then we note that Nash equilibria always exist in these systems as well. If we remove the finite-horizon specification, then the status of the realizability problem for infinite-horizon systems now becomes a long-standing open problem in the field of game theory. In [CMJ04], it was claimed that there existed a 2-agent game infinite-horizon game in which each of the agents had a reachability goal that did not have a Nash equilibrium. This claim, however, was shown to be flawed in [BMS14], so the question of whether the realizability problem is trivial in this setting is still open. Since even this is not known, the complexity of the realizability problem for Nash equilibria remains open in the infinite-horizon setting. Similarly, it is not known if a subgame perfect equilibrium strategy profile always exists in the same unbounded setting.
@@
This leads us to direct our focus on the verification problem for both the Nash and subgame perfect equilibrium concepts. The verification problem takes an input of both a system and a candidate tuple of strategies and determines whether the input tuple of strategies satisfies the equilibrium condition. In this way, it can naturally be seen as dual to the often considered problem of computing an equilibrium from scratch that is foundational to fields like algorithmic game theory [Rou10]. However, despite the fact that recent works [RV22, Umm11, Par23, KNPS21] and tools [KNPS20, GNPW18, CGSP15] have begun to consider practical aspects of the verification problem, there has been virtually no theoretical development of the verification problem in probabilistic settings.
@@
In order to accurately isolate the complexity of verification, we introduce a new succinctly represented model in this paper, the b-bounded concurrent game system. In a b-bounded system, only b agents select an action concurrently at all given game-states, where b is a fixed constant. This model serves two purposes. First, it allows us to generalize from the popular but restrictive turn-based game model that is ubiquitous in the literature [Umm11, Gup16, Woo09, Rou10]. Second, it sharpens our complexity-theoretic results by circumventing the construction of an exponentially large transition table as part of our input. In a multi-agent system, the transition table of an unrestricted model can be seen as an exponential construction in itself, a point we discuss explicitly in Section 2 and that has been noted in other works such as [BBMU15]. This exponential construction within the input devalues the complexity-theoretic value of upper bounds, as a PSPACE upper bound would practically translate into doubly-exponential time complexity in the number of agents. We discuss how our results extend to a full concurrency model in Section 3.6 (one in which all agents select an action at all game-states). Furthermore, the issue of goal representation also plays a significant role in the overall computation complexity of the verification problem. As noted in [RV21, RV22], if agents are given highly succinct temporal logic formulae that represent their preferences on the system’s executions, then the worst-case complexity of reasoning about these logics can completely overpower the worst-case complexity of the actual verification decision problem. In order to avoid this problem, we give the agents reachability goals, which are arguably the simplest non-trivial goals.
@@
Our analysis starts with the subgame perfect equilibrium. For the subgame perfect equilibrium verification problem, we are able to show a PSPACE upper bound by employing results from circuit complexity and carefully considering issues of numerical representation that are often overlooked in the literature. As we demonstrate later on, a key subroutine in the verification decision problem is solving an exponentially-large system of equations. In [Put94](Theorem 3.2, Corollary 3.2b), it is shown that systems of equations with rational coefficients admit rational solutions with polynomial-size encodings, meaning that for an exponentially-large system of equations, we would get an exponentially-large representation of a rational number solution in the worst-case. In this paper, we consider a numerical model directly inspired by the finite-precision arithmetic models used in real programs and explicitly construct a bound on solution sizes that is unique to our setting and model. These bounds allow us to get a PSPACE upper bound for the subgame perfect equilibrium verification problem. If we note that backwards induction would compute a subgame perfect equilibrium from scratch in EXPTIME, then this result lines up with our intuition, as pending a PSPACE=EXPTIME result, there is a computational benefit to verifying a given set of input strategies as opposed to explicitly synthesizing them.
@@ 2. Preliminaries
We start by introducing the main object of our study, b-bounded concurrent game system. This model is very similar to the well-studied “multi-player concurrent stochastic game” model [Sol22]. The major difference between our model and the most standard concurrent game models in the literature is the “b-bounded” property.
@@
    $\delta$ is the transition function. Let $v,v^{\prime}$ be elements of $V$ and let $\theta$ be an element of $\Theta_{P(v)}$. The transition function $\delta$ maps triples $(v,\theta,v^{\prime})$ to $[0,1]$. These outputs represent a probability distribution, and we therefore have $\sum_{v^{\prime}\in V}\delta(v,\theta,v^{\prime})=1$. Intuitively, the transition function only considers the actions of agents that are included in the image of the playing function when considering possible transitions from a state. Note that it is possible that $P(v)=\emptyset$. In this case, we say that $v$ is uncontrolled as the transitions from state $v$ are determined by a fixed probability distribution that does not take input from the agents, analogous to uncontrollable “probabilistic” states in the theory of Markov Decision Processes [Put94].
@@
Note that in this definition, we have $1$-indexed the set of agents as opposed to 0-indexing that is more present in the related computer science literature. This is because the model presented more closely lines up with models studied in game theory [Sol22, FT91], and the game theory literature adopts mathematical conventions such as $1$-indexing as opposed to computer science conventions such as 0-indexing.
Taken together, our concurrent game model represents a very well-studied standard finite-horizon stochastic game model [Sol22] with one exception: the aforementioned b-bounded property. In this model, we have a bound on the concurrency of the system given by the playing function $P$ and the bound $b$. Note that in order for this model to be considered a “game” where agents make decisions, we must have $b\geq 1$. This bound was introduced for the purpose of complexity-theoretic analysis, as, without this bound, the transition table $\delta$ can be seen as an exponential construction itself. For example, if we have $k$ agents and each agent has two actions, a transition table would need to have $|V|^{2}\times 2^{k}$ entries in the worst-case scenario. This table would represent the probability that a state $v_{1}$ transitions to a second state $v_{2}$ under each of the $2^{k}$ collective action choices in the full action product space $\Theta_{\Omega}$. The construction of such a large transition table would devalue the meaning of the complexity-theoretic results presented in this paper, as an exponential time result in the size of the game input would be interpreted as doubly-exponential in the number of agents. This led us to choose a natural restriction of concurrent games that would have a transition table with a more succinct presentation. In a b-bounded game, the transition table has at most $|V|^{2}\times 2^{b}$ entries for a fixed $b$ (and therefore polynomial in the number of agents). This allows us to demonstrate upper bounds that hold even for models that have a more compact transition table representation. Furthermore, the new size bound allows us to better tie our complexity results to the number of agents and states in the game, as the transition table no longer adds an extra exponential blowup to our complexity-theoretic analysis. In Section 3.6, we explicitly consider the upper bound for a full concurrency model and show that the PSPACE membership result of Theorem 7 still holds for this model as well.
@@
In our model, probabilistic transducers have a deterministic transition function and a finite state space. A discussion of why considering such transducers is sufficient and what would happen if a probabilistic transition function was used instead can be found in Section 3.6. Since transducers are provided as part of our input, we once again assume that all probabilities output by $O_{i}$ are exactly $\mathbb{L}$ (by slight abuse of notation, $\mathbb{L}$ was defined in $\mathbb{G}$) bits in length to deal with issues of numerical representation. Unlike the deterministic transition function and the finite state space, this is a restriction on the model as it does not allow us to represent repeating decimals or irrational numbers. For repeating decimals, we note that the numbers do not necessarily have to be represented in base two, though we assume this for our computations. For irrational numbers, there exist games in which the only equilibria involve distributions with irrational probabilities [Nas51]. Nevertheless, the verification problem is ultimately a problem about programs, and real programs can only work with finite precision, so a finite representation of probabilities is a reasonable restriction. A further extension of this work would be to consider rational-number representations of probabilities, which can be seen as another “finite” representation. This would introduce new issues with numerical representation about how to represent and perform operations on the rational representations of the probabilities.
@@ 3. Subgame Perfect Equilibrium Upper Bound
Given two states $s,r\in\mathbb{G}\times\pi$, it is possible to compute the transition probability $P(s,r)$ in PSPACE w.r.t to the input of $\mathbb{G}$ and $\pi$
@@ 3.1. Outline
Section Goal
Section 3.2 Establish a correspondence between subgames in $\mathbb{G}$ and a set of states $R$ in $\mathbb{G}\times\pi$
Section 3.3 Establish a PSPACE algorithm for computing hitting probabilities in $\mathbb{G}\times\pi$
Section 3.4 Prove that if $\pi$ is not a subgame perfect equilibrium, this can be witnessed by a “small, local” change at a state $r\in R$ that results in a strictly higher hitting probability at $r$
Section 3.5 Compare the two hitting probabilities from Section 3.3 (no deviation) and Section 3.4 (after a “small” improvement) from the state $r\in R$ in PSPACE
Section 3.6 Consider alternate models
Figure 1. A description of the sections to follow.
@@ 3.3. Computing Winning Probabilities From a State
Following our outline, the next step is to compute the probability that Agent $i$ reaches his goal in $\mathbb{G}|_{h}$ when all agents (including $i$) follow $\pi|_{h}$ for some history $h$. Lemma 2 gives us a way to guess an arbitrary history in PSPACE. Therefore, we assume that both the Agent $i$ and the history $h$ are fixed in this section and denote the relevant history-reachable state resulting from $h$ in $\mathbb{G}\times\pi$ as $r$. Starting from the state $r\in R$ in $\mathbb{G}\times\pi$, we want to compute the probability that a member of the set $T=\{\langle v,s,n\rangle|v\in G_{i}\}$ is reached, as this is equivalent to the probability that Agent $i$ reaches a state in his goal set in $\mathbb{G}|_{h}$ when the strategy $\pi|_{h}$ is followed. In the Markov Chain $\mathbb{G}\times\pi$, the probability that a member of some set is visited when starting from some state is called a hitting probability [Put94], and so we are interested in the hitting probability $h_{r}$ of $T$ when started in state $r$.
A standard way to solve for a hitting probability is to create a system of linear equations [Put94]. In $\mathbb{G}\times\pi$, for states $t\in T$, we would use the equation $h_{t}=1$ – a state already in $T$ visits $T$ with probability $1$. For other states $r$, we would use the equation $h_{r}=\sum_{v\in V\times S\times[F]}P(r,v)\cdot h_{v}$ – this relates the hitting probability of $r$ to the hitting probabilities of the states its transitions to. Taken together, the hitting probabilities would be computed by a set of $|V\times S\times[F]|$ equations over $|V\times S\times[F]|$ states. However, it is possible for this system to be underdetermined, meaning that standard linear algebra techniques would not apply.
@@
Consider a Markov Chain $\mathcal{C}$ with two non-target states $x$ and $y$ with the following properties : The state $x$ self transitions with probability $\frac{1}{2}$ and transitions to $y$ with probability $\frac{1}{2}$. Furthermore, the state $y$ self transitions with probability $\frac{1}{2}$ and transitions to $x$ with probability $\frac{1}{2}$ – see Figure 2 for an illustration. The equations for the hitting probabilities $h_{x}$ and $h_{y}$ would be both given by the linear equation $\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, meaning that the system of equations used to solve for hitting probabilities in $\mathcal{C}$ would be underdetermined and have infinite solutions.
@@
Figure 2. An illustration of Example $3.$ The entire figure represents the Markov Chain $\mathcal{C}$, which has two disconnected components: $\mathcal{C^{\prime}}$ and the two non-target states $x$ and $y$. We wish to calculate the hitting probabilities for some state $c\in\mathcal{C^{\prime}}$ by setting up a system of linear equations. If we set the hitting probabilities up in a naïve way, we get $h_{x}=h_{y}=\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, a pair of equations with an infinite number of solutions. Thus, the system of equations to solve for the hitting probabilities in $\mathcal{C}$ is underdetermined.
@@
Solving a system of explicitly given linear equations with a unique solution belongs to the class NC2 [Csa76], a subset of the class NC of problems solvable in parallel poly-logarithmic time using polynomially many processors [Coo85]. The standard solution concept for a system of linear equations is a vector $\bf v$} that associates a numerical value to each variable $v$ in the system of equations. Since NC2 is a complexity class containing decision problems [Joh90], the algorithm in [Csa76] technically takes as input a variable $v$ and an index $k$ and decides the value of $k$-th bit of $v$ in $\bf v$}. The class NC2 itself contained in log2-SPACE [Joh90], meaning that we can create a SPACE($\log^{2}n)$ Turing Machine $\mathcal{T}$ that is capable of simulating the NC2 algorithm presented in [Csa76]. The machine $\mathcal{T}$ takes a variable $v$, and an index $k$ as input, and it simulates the relevant processors of the NC2 algorithm to decide whether the $k$-th bit of the value of $v$ in the solution vector $\bf v$} is $1$ or 0. Since $\mathcal{T}$ runs in poly-logarithmic space, $\mathcal{T}$ would have a read-only input tape that explicitly describes the system of equations, the variable $v$, and the index $k$, alongside an additional separate work tape that uses $O(\log^{2}n)$ space [Sip06].

### v6

@@ Abstract.
Finite-horizon probabilistic multiagent concurrent game systems, also known as finite multiplayer stochastic games, are a well-studied model in computer science due to their ability to represent a wide range of real-world scenarios involving strategic interactions among agents over a finite amount of iterations (given by the finite-horizon). The analysis of these games typically focuses on evaluating (verifying) and computing (synthesizing/realizing) which strategy profiles (functions that represent the behavior of each agent) qualify as equilibria. The two most prominent equilibrium concepts are the Nash equilibrium and the subgame perfect equilibrium, with the latter considered a conceptual refinement of the former. Computing these equilibria from scratch is, however, often computationally infeasible. Therefore, recent attention has shifted to the verification problem, where a given strategy profile must be evaluated to determine whether it satisfies equilibrium conditions. In this paper, we demonstrate that the verification problem for subgame perfect equilibria lies in PSPACE, while for Nash equilibria, it is EXPTIME-complete. This is a highly counterintuitive result since subgame perfect equilibria are often seen as a strict strengthening of Nash equilibria and are intuitively seen as more complicated.
@@ 1. Introduction
The ubiquity of distributed systems in modern computing architectures has motivated much recent work into the study of multi-agent systems [18, 19, 27, 40, 44, 45, 1, 35, 37, 36, 6, 4, 3, 5, 8, 7], which are abstract mathematical models in which two or more rational agents interact with each other in order to achieve their goals. The broad motivation behind these works is to provide a mathematical framework for analyzing the overall trends of the complex mechanics that arise when multiple agents interact with one another in a strategic manner.
@@
Much of the relevant work in this area comes from the field of formal methods, which has recently begun to analyze equilibria in multi-agent systems as a way to analyze “stable” behaviors in multi-agent systems, and, therefore, analyze the system’s properties as a whole. The two major problems in the field of formal methods are the realizability [33] and the verification problems [38, 37]. Therefore, much of the work in formal analysis of multi-agent systems is concerned with determining the existence of (the realizability problem) and verifying (the verification problem) different types of equilibria [18, 4, 6, 5, 35, 37, 36, 19, 21, 22, 5, 8, 7]. As evidenced by the previously cited works, however, work in formal verification often comes with a strong preference for deterministic systems and deterministic agent behaviors. This creates something of an ideological divide between equilibrium analysis in the formal methods literature and the study of equilibria in the game theory literature. Our broad goal in this paper is to begin to bridge this divide by applying this formal methods style of equilibrium analysis to probabilistic systems. This is a new line of inquiry, as even papers in the formal methods literature that consider probabilistic systems such as [20] (which itself notes the focus on deterministic systems in previous literature) consider notions of almost-sure satisfaction on infinite-length executions of a system. These notions allow for these probabilistic systems to be analyzed through the same techniques used for deterministic ones.
@@
In order to differentiate our work from this type of analysis, we consider finite-horizon systems, meaning that the system has an explicitly given finite bound on the number of steps in the execution of the system. This terminology follows the notation from the literature on Markov Decision Processes [34, 14] and breaks from the terminology of the formal verification literature, where finite-horizon is usually shorthand for “finite but unbounded” [45, 17]. Finite-horizon systems are also arguably more popular than their infinite-horizon counterparts in the literature on Markov Decision Process [34, 14, 42, 2], which can be seen as the “single-agent” analog of the systems we consider in this paper.
@@
We wish to consider the computational complexity of the realizability and verification problems in finite-horizon probabilistic multi-agent systems. The two most popular equilibrium concepts in the field of game theory are the Nash equilibrium [28] and the subgame perfect equilibrium [30], with the latter a strengthening of the former. However, this immediately presents a problem for the realizability problem. Since the realizability problem is concerned with determining whether an equilibrium exists and we are considering probabilistic systems with a finite-horizon, the realizability problem becomes trivial. This is because subgame perfect equilibria can be shown to always exist in these systems through the use of algorithms like backwards induction [15]. Furthermore, if we note that every subgame perfect equilibrium is, by definition, a Nash equilibrium, then we note that Nash equilibria always exist in these systems as well. If we remove the finite-horizon specification, then the status of the realizability problem for infinite-horizon systems now becomes a long-standing open problem in the field of game theory. In [11], it was claimed that there existed a 2-agent game infinite-horizon game in which each of the agents had a reachability goal that did not have a Nash equilibrium. This claim, however, was shown to be flawed in [3], so the question of whether the realizability problem is trivial in this setting is still open. Since even this is not known, the complexity of the realizability problem for Nash equilibria remains open in the infinite-horizon setting. Similarly, it is not known if a subgame perfect equilibrium strategy profile always exists in the same unbounded setting.
@@
This leads us to direct our focus on the verification problem for both the Nash and subgame perfect equilibrium concepts. The verification problem takes an input of both a system and a candidate tuple of strategies and determines whether the input tuple of strategies satisfies the equilibrium condition. In this way, it can naturally be seen as dual to the often considered problem of computing an equilibrium from scratch that is foundational to fields like algorithmic game theory [39]. However, despite the fact that recent works [37, 43, 32, 25] and tools [24, 21, 9] have begun to consider practical aspects of the verification problem, there has been virtually no theoretical development of the verification problem in probabilistic settings.
@@
In order to accurately isolate the complexity of verification, we introduce a new succinctly represented model in this paper, the b-bounded concurrent game system. In a b-bounded system, only b agents select an action concurrently at all given game-states, where b is a fixed constant. This model serves two purposes. First, it allows us to generalize from the popular but restrictive turn-based game model that is ubiquitous in the literature [43, 19, 45, 39]. Second, it sharpens our complexity-theoretic results by circumventing the construction of an exponentially large transition table as part of our input. In a multi-agent system, the transition table of an unrestricted model can be seen as an exponential construction in itself, a point we discuss explicitly in Section 2 and that has been noted in other works such as [5]. This exponential construction within the input devalues the complexity-theoretic value of upper bounds, as a PSPACE upper bound would practically translate into doubly-exponential time complexity in the number of agents. We discuss how our results extend to a full concurrency model in Section 3.6 (one in which all agents select an action at all game-states). Furthermore, the issue of goal representation also plays a significant role in the overall computation complexity of the verification problem. As noted in [36, 37], if agents are given highly succinct temporal logic formulae that represent their preferences on the system’s executions, then the worst-case complexity of reasoning about these logics can completely overpower the worst-case complexity of the actual verification decision problem. In order to avoid this problem, we give the agents reachability goals, which are arguably the simplest non-trivial goals.
@@
Our analysis starts with the subgame perfect equilibrium. For the subgame perfect equilibrium verification problem, we are able to show a PSPACE upper bound by employing results from circuit complexity and carefully considering issues of numerical representation that are often overlooked in the literature. As we demonstrate later on, a key subroutine in the verification decision problem is solving an exponentially-large system of equations. In [34](Theorem 3.2, Corollary 3.2b), it is shown that systems of equations with rational coefficients admit rational solutions with polynomial-size encodings, meaning that for an exponentially-large system of equations, we would get an exponentially-large representation of a rational number solution in the worst-case. In this paper, we consider a numerical model directly inspired by the finite-precision arithmetic models used in real programs and explicitly construct a bound on solution sizes that is unique to our setting and model. These bounds allow us to get a PSPACE upper bound for the subgame perfect equilibrium verification problem. If we note that backwards induction would compute a subgame perfect equilibrium from scratch in EXPTIME, then this result lines up with our intuition, as pending a PSPACE=EXPTIME result, there is a computational benefit to verifying a given set of input strategies as opposed to explicitly synthesizing them.
@@ 2. Preliminaries
We start by introducing the main object of our study, b-bounded concurrent game system. This model is very similar to the well-studied “multi-player concurrent stochastic game” model [42]. The major difference between our model and the most standard concurrent game models in the literature is the “b-bounded” property.
@@
    $\delta$ is the transition function. Let $v,v^{\prime}$ be elements of $V$ and let $\theta$ be an element of $\Theta_{P(v)}$. The transition function $\delta$ maps triples $(v,\theta,v^{\prime})$ to $[0,1]$. These outputs represent a probability distribution, and we therefore have $\sum_{v^{\prime}\in V}\delta(v,\theta,v^{\prime})=1$. Intuitively, the transition function only considers the actions of agents that are included in the image of the playing function when considering possible transitions from a state. Note that it is possible that $P(v)=\emptyset$. In this case, we say that $v$ is uncontrolled as the transitions from state $v$ are determined by a fixed probability distribution that does not take input from the agents, analogous to uncontrollable “probabilistic” states in the theory of Markov Decision Processes [34].
@@
Note that in this definition, we have $1$-indexed the set of agents as opposed to 0-indexing that is more present in the related computer science literature. This is because the model presented more closely lines up with models studied in game theory [42, 16], and the game theory literature adopts mathematical conventions such as $1$-indexing as opposed to computer science conventions such as 0-indexing.
Taken together, our concurrent game model represents a very well-studied standard finite-horizon stochastic game model [42] with one exception: the aforementioned b-bounded property. In this model, we have a bound on the concurrency of the system given by the playing function $P$ and the bound $b$. Note that in order for this model to be considered a “game” where agents make decisions, we must have $b\geq 1$. This bound was introduced for the purpose of complexity-theoretic analysis, as, without this bound, the transition table $\delta$ can be seen as an exponential construction itself. For example, if we have $k$ agents and each agent has two actions, a transition table would need to have $|V|^{2}\times 2^{k}$ entries in the worst-case scenario. This table would represent the probability that a state $v_{1}$ transitions to a second state $v_{2}$ under each of the $2^{k}$ collective action choices in the full action product space $\Theta_{\Omega}$. The construction of such a large transition table would devalue the meaning of the complexity-theoretic results presented in this paper, as an exponential time result in the size of the game input would be interpreted as doubly-exponential in the number of agents. This led us to choose a natural restriction of concurrent games that would have a transition table with a more succinct presentation. In a b-bounded game, the transition table has at most $|V|^{2}\times 2^{b}$ entries for a fixed $b$ (and therefore polynomial in the number of agents). This allows us to demonstrate upper bounds that hold even for models that have a more compact transition table representation. Furthermore, the new size bound allows us to better tie our complexity results to the number of agents and states in the game, as the transition table no longer adds an extra exponential blowup to our complexity-theoretic analysis. In Section 3.6, we explicitly consider the upper bound for a full concurrency model and show that the PSPACE membership result of Theorem 7 still holds for this model as well.
@@
In our model, probabilistic transducers have a deterministic transition function and a finite state space. A discussion of why considering such transducers is sufficient and what would happen if a probabilistic transition function was used instead can be found in Section 3.6. Since transducers are provided as part of our input, we once again assume that all probabilities output by $O_{i}$ are exactly $\mathbb{L}$ (by slight abuse of notation, $\mathbb{L}$ was defined in $\mathbb{G}$) bits in length to deal with issues of numerical representation. Unlike the deterministic transition function and the finite state space, this is a restriction on the model as it does not allow us to represent repeating decimals or irrational numbers. For repeating decimals, we note that the numbers do not necessarily have to be represented in base two, though we assume this for our computations. For irrational numbers, there exist games in which the only equilibria involve distributions with irrational probabilities [29]. Nevertheless, the verification problem is ultimately a problem about programs, and real programs can only work with finite precision, so a finite representation of probabilities is a reasonable restriction. A further extension of this work would be to consider rational-number representations of probabilities, which can be seen as another “finite” representation. This would introduce new issues with numerical representation about how to represent and perform operations on the rational representations of the probabilities.
@@ 3. Subgame Perfect Equilibrium Upper Bound
Given two states $s,r\in\mathbb{G}\times\pi$, it is possible to compute the transition probability $P(s,r)$ in PSPACE w.r.t to the size of the input of $\mathbb{G}$ and $\pi$
@@ 3.1. Outline
Figure 1. A chart depiction of the outline with some figures intended to facilitate conceptual understanding.
@@ 3.3. Computing Winning Probabilities From a State
Following our outline, the next step is to compute the probability that Agent $i$ reaches his goal in $\mathbb{G}|_{h}$ when all agents (including $i$) follow $\pi|_{h}$ for some history $h$. Lemma 2 gives us a way to guess an arbitrary history in PSPACE. Therefore, we assume that both the Agent $i$ and the history $h$ are fixed in this section and denote the relevant history-reachable state resulting from $h$ in $\mathbb{G}\times\pi$ as $r$. Starting from the state $r\in R$ in $\mathbb{G}\times\pi$, we want to compute the probability that a member of the set $T=\{\langle v,s,n\rangle|v\in G_{i}\}$ is reached, as this is equivalent to the probability that Agent $i$ reaches a state in his goal set in $\mathbb{G}|_{h}$ when the strategy $\pi|_{h}$ is followed. In the Markov Chain $\mathbb{G}\times\pi$, the probability that a member of some set is visited when starting from some state is called a hitting probability [34], and so we are interested in the hitting probability $h_{r}$ of $T$ when started in state $r$.
A standard way to solve for a hitting probability is to create a system of linear equations [34]. In $\mathbb{G}\times\pi$, for states $t\in T$, we would use the equation $h_{t}=1$ – a state already in $T$ visits $T$ with probability $1$. For other states $r$, we would use the equation $h_{r}=\sum_{v\in V\times S\times[F]}P(r,v)\cdot h_{v}$ – this relates the hitting probability of $r$ to the hitting probabilities of the states its transitions to. Taken together, the hitting probabilities would be computed by a set of $|V\times S\times[F]|$ equations over $|V\times S\times[F]|$ states. However, it is possible for this system to be underdetermined, meaning that standard linear algebra techniques would not apply. To demonstrate this, we abstract away from our setting and illustrate how a standard linear-algebraic approach could go wrong in Example 3.3. This example is intended to inform readers unfamiliar with the application of linear-algebraic techniques to Markov Chains of the general setting; we then discuss why this potential error does not apply to our setting.
@@
Consider a Markov Chain $\mathcal{C}$ with two non-target states $x$ and $y$ with the following properties: The state $x$ self-transitions with probability $\frac{1}{2}$ and transitions to $y$ with probability $\frac{1}{2}$. Furthermore, the state $y$ self transitions with probability $\frac{1}{2}$ and transitions to $x$ with probability $\frac{1}{2}$ – see Figure 2 for an illustration. The equations for the hitting probabilities $h_{x}$ and $h_{y}$ would be both given by the linear equation $\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, meaning that the system of equations used to solve for hitting probabilities in $\mathcal{C}$ would be underdetermined and have infinite solutions.
@@
Figure 2. An illustration of Example 3.3. The entire figure represents the Markov Chain $\mathcal{C}$, which has two disconnected components: $\mathcal{C^{\prime}}$ and the two non-target states $x$ and $y$. We wish to calculate the hitting probabilities for some state $c\in\mathcal{C^{\prime}}$ by setting up a system of linear equations. If we set the hitting probabilities up in a naïve way, we get $h_{x}=h_{y}=\frac{1}{2}h_{x}+\frac{1}{2}h_{y}$, a pair of equations with an infinite number of solutions. Thus, the system of equations to solve for the hitting probabilities in $\mathcal{C}$ is underdetermined.
@@
Solving a system of explicitly given linear equations with a unique solution belongs to the class NC2 [13], a subset of the class NC of problems solvable in parallel poly-logarithmic time using polynomially many processors [12]. The standard solution concept for a system of linear equations is a vector $\bf v$} that associates a numerical value to each variable $v$ in the system of equations. Since NC2 is a complexity class containing decision problems [23], the algorithm in [13] technically takes as input a variable $v$ and an index $k$ and decides the value of $k$-th bit of $v$ in $\bf v$}. The class NC2 itself contained in log2-SPACE [23], meaning that we can create a SPACE($\log^{2}n)$ Turing Machine $\mathcal{T}$ that is capable of simulating the NC2 algorithm presented in [13]. The machine $\mathcal{T}$ takes a variable $v$, and an index $k$ as input, and it simulates the relevant processors of the NC2 algorithm to decide whether the $k$-th bit of the value of $v$ in the solution vector $\bf v$} is $1$ or 0. Since $\mathcal{T}$ runs in poly-logarithmic space, $\mathcal{T}$ would have a read-only input tape that explicitly describes the system of equations, the variable $v$, and the index $k$, alongside an additional separate work tape that uses $O(\log^{2}n)$ space [41].